A monopoly firm has the following demand curve: Q = 2,000 – 25P
where Q is its monthly output. Assuming its monthly short-run total cost is described by the function
STC = 500 + 8Q + 0.035Q2 Answer the following questions:
i. What will be its profit-maximising price and output?
ii. How much profit will it have at the preceding output?
1
Expert's answer
2020-12-23T11:00:48-0500
"p=80-0.04Q"

"TR=80Q-0.04Q^2"


"MR=80-0.08Q"


"MC=8+0.07Q"


"MR=MC"


"Q=480"


"p=60.8"


"\pi=TR-STC=72Q-0.07Q^2-500"


"\pi=17,932"